# Channel Choice and Customer Value

### v1

@@ 2 Related Literature
Moreover, the setting of our study is a fast-moving consumer packaged goods retailer (pet supplies), whereas much of the earlier literature uses different settings like books, apparel, video games, electronics, technology, etc. These earlier settings differ from ours on many dimensions, e.g., they tend to be less frequently purchased, do not have as many competing stores, and the nature of the product is more hedonic (rather than utilitarian), which has been found to be important for the multichannel effect (Kushwaha and Shankar 2013).
Most importantly, the existing literature overlooks the heterogeneity in channel adoption pathways, implicitly assuming that the benefits derived from multichannel customers are consistent, irrespective of the motivations underlying their adoption decisions. However, theoretical frameworks suggest important distinctions between (a) organic adopters and those induced through external incentives, and (b) different types of external stimuli—namely, macro-environmental shocks versus firm-initiated promotions. Consumer inertia and habit theory (Verplanken and Wood 2006; Wood and Neal 2009) suggests that externally triggered adoption events—such as those prompted by environmental disruptions or marketing campaigns—may temporarily disrupt existing behavioral routines, resulting in short-term channel experimentation. Without reinforcement, however, such consumers are likely to revert to prior habits, thereby potentially limiting the long-term impact of multi-channel adoption. Moreover, when adoption is driven by promotions, the literature on forward buying and demand distortion (Neslin, Henderson, and Quelch 1985; Ailawadi and Neslin 1998) raises concerns about post-promotion gains for firms. Specifically, promotion-induced online channel adoption may lead to intertemporal demand shifts rather than incremental sales, reducing post-promotion revenues and potentially diminishing the value of these new multichannel users.
Thus, this paper extends the literature from a substantive perspective in two key ways: First, we focus on a key type of channel transition: when customers move from shopping only in offline stores to using both offline and online channels. Second, we examine how post-adoption behavior differs depending on why consumers started using the online channel— whether organically or due to external factors like promotions or macro-environmental shocks, and provide theoretical explanations for the differences. By doing so, we challenge the implicit assumption in prior research that all multichannel customers behave similarly post-adoption and are equally valuable to firms, regardless of how or why they adopted a new channel [^1] Like us, Montaguti, Neslin, and Valentini (2016) also consider an external stimulus for multichannel adoption (marketing campaign). However, the methods, goals, and estimands of interest of the two papers are quite different – we use difference-in-differences to study the change caused by adoption of online shopping, whereas they use a cross-sectional comparison to measure the effectiveness of a marketing campaign nudging consumers to use multiple channels. The substantive settings also differ – we focus on the transition from using only physical stores to physical stores + e-commerce in a competitive pet groceries setting, whereas they use data from a bookstore with a subscription model and focus on subscribers who use single or multiple channels (mail order, fax, phone, SMS, the Internet, and bookstores). Most importantly, they do not study the differences between organic multichannel customers vs. those who were nudged by the marketing campaign, which is a key area of focus in our paper..
Next, from a methodological perspective, prior research has considered three broad approaches to documenting the value of adopting new channels: cross-sectional, within-user, and DID. Below, we discuss each stream and explain how our research is distinct from the earlier work.
Cross-sectional: Papers in this stream (Thomas and Sullivan 2005; Kumar and Venkatesan 2005; Kushwaha and Shankar 2013; Montaguti, Neslin, and Valentini 2016) typically compare the revenue/profits of two distinct groups of consumers to document the value of multichannel shopping – those who use multiple channels (among catalog, brick-and-mortar, salespersons, direct mail, telephone sales, SMS or the internet) vs. those who use a single channel. A consensus finding in this literature is that multichannel customers tend to be more valuable than single-channel customers. However, a potential drawback of this methodology is that multichannel consumers are self-selected and could be systematically different from those using single channels. As such, cross-sectional comparisons cannot answer the question of how customer behavior changes after the adoption of multiple channels.
Within Customer: Papers in this stream (Venkatesan and Ravishanker 2007; Ansari, Mela, and Neslin 2008) typically compare how a customer’s purchase behavior changes after the adoption of multiple channels (which could be different types of stores, catalogs, internet etc.) compared to their behavior when they were a single-channel customer. The key takeaway here is that customers spend more after becoming multichannel compared to when they were using only a single channel. However, there are important drawbacks to this type of analysis. First, without an appropriate control group of non-adopters, we need to assume that the entire difference between the before and after periods can be fully attributed to the adoption of multiple channels, and there are no other time-varying factors affecting purchasing behavior – which is a strong assumption. Second, it is not obvious if the estimated gains of the adopters can be extrapolated to non-adopters because the adopters self-selected into the use of multiple channels.
Difference-in-Differences (DID): This more recent stream of literature tries to address some of the concerns with the above two approaches by using a DID methodology and using non-adopters as a control group. However, the papers in this stream (Wang, Malthouse, and Krishnamurthi 2015; Narang and Shankar 2019; Kim, Wang, and Malthouse 2015; Gu and Kannan 2021) focus only on the adoption of mobile apps rather than a broader set of channels. Specifically, they find that customers who adopted a firm’s mobile app (from previously using physical stores, website, or both) increase their purchases post-adoption [^2] Gu and Kannan (2021) shows the opposite result, i.e., a decrease in purchases post-app adoption for a hotel chain. While our paper also uses a DID method, our substantive focus is very different from these papers, as explained previously.
Finally, while the literature on multichannel vs. single-channel shopping described above is most closely related to our paper, there are a couple of other strands of the marketing literature that are also somewhat related. First, there is a literature that has studied the role of channels in non-retail settings, e.g., quantifying the effects of the adoption of new channels (like streaming services) on the downloading and owning in the context of entertainment goods like movies and music (Datta, Knox, and Bronnenberg 2018; Aguiar and Waldfogel 2018). Second, there is also literature that studies related questions like the value of physical stores in an omnichannel environment (Zhang, Chang, and Neslin 2022) and the effects of customer experiences in different channels on retention (Chang and Zhang 2016). See Zhang et al. (2010) and Cui et al. (2021) for a summary of the findings and challenges in the broader literature on omnichannel retail strategy.
@@ 3.1 Setting
Our data come from one of the largest Brazilian pet supplies retailers, which sells a wide variety of products for pets, such as pet food and snacks, medicines/pharmacy items, hygiene and grooming products, and various pet toys and accessories. The firm has both offline and online channels. The retailer operates more than 200 physical stores in Brazil, which serve as the offline channel, and has a website as well as a mobile app, which serve as the online channel. Pet supplies is a large and competitive industry with about \$250 billion in annual revenue globally (Fortune 2024).
@@ 3.2 Data
Our data include transaction-level records of all individual customers from both online and offline channels during the 63-month period from January 1st, 2019 to March 31st, 2024. Each transaction record is uniquely identified by an order number with one or more purchased items. For each purchased item, we have information on the unit price, quantity purchased, brand, purchased channel (online/offline), subcategory, and category (e.g., food, hygiene).
@@
A potential concern with any channel analysis is that the results could be driven by assortment differences across the two channels. To ensure that this is not the case, we exclude data from four product categories that were not available in both channels for the entirety of the observation period; see Table A1 in Web Appendix $\S$§A for a summary of these categories and their availability information. Together, sales from these categories account for only 0.63% of the total sales during the observation period; as such we do not expect these to play any significant role in the findings.
The total number of consumers who shopped at least once during our sample period is 5,934,248444This is the set of consumers who purchased at least one item from one of the non-excluded categories.. For each of these consumers, we aggregate the transaction-level data into customer-month-level panel data over the customer’s entire purchase history within the observation period (January 2019 to March 2024), starting from his or her initial purchase month and ending with the last observed purchase month. For example, if a customer purchased for the first time in our data in August 2019 and the last observed purchase is in July 2020, then the panel for this customer spans 12 months, from August 2019 to July 2020. If there are intermediate months in which a customer did not make any purchase, the values for the monthly purchase-related variables are set to zero.
We now discuss the summary statistics of our data. In $\S$3.2.1, we summarize the consumer-level demographic data, and in $\S$3.2.2, we summarize the monthly purchase behaviors.
3.2.1 Time Invariant Consumer Attributes
Table 2: Summary Statistics of Customer-Level Categorical Demographic Variables
Gender Count Percentage (%)
Female 2,351,011 39.62
Male 1,521,157 25.63
No Information 2,062,080 34.75
Total 5,934,248 100.00
Table 3: Summary Statistics of Customer-Level Demographic Variables
Variable Mean SD 25th 50th 75th (Min, Max) Count Missing (%)
Age 40.96 12.75 31 39 49 (7, 99) 4,302,771 27.49
HouseholdIncome 1854.37 1471.55 824.54 1358.21 2448.73 (8, 30826) 4,377,957 26.23
PanelLength 19.76 19.39 1 14 34 (1, 63) 5,934,248 0.00
We first summarize the time-invariant attributes of the consumers in the data. We have data on five consumer-specific variables:
  • $\bullet$
    $Gender_{i}$: Categorical variable denoting $i^{\prime}s$ gender (Male, Female, or No Information).
  • $\bullet$
    $Age_{i}$: Customer $i$’s age.
  • $\bullet$
    $HouseholdIncome_{i}$: Customer $i^{\prime}s$ average monthly household income.
  • $\bullet$
    $PanelLength_{i}$: Period spanning the first and last purchases for customer $i$ (in months).
  • $\bullet$
    $Zipcode_{i}$: The zipcode of customer $i$.
We present the summary statistics for demographic variables in Tables 2 and 3. Out of the 5,934,248 customers, 2,351,011 (39.6%) are Female, 1,521,157 (25.6%) are Male, and the rest (34.8%) do not report gender information. The median consumer is 39 years old. In addition, the mean and median of monthly household incomes are 1854.37 and 1358.21 Brazilian Real, respectively. Next, we see that the average and median panel lengths are 19.76 and 14 months respectively, which suggests that there is consumer entry as well as churn during the observation period. Finally, we observe that customers in our data sample come from 300,774 distinct zip codes across all of the 28 states of Brazil (though 20.4% of total customers do not provide the zip code information).
3.2.2 Customer Monthly Purchase Behaviors
Table 4: Customer-Month Summary Statistics
Variables mean std min 25% 50% 75% max count
Transactions across both Channels
Spend 303.02 759.25 0 0 0 374.92 2,340,617.16 123,748,452
Quantity 2.44 10.27 0 0 0 2.00 44,158.00 123,748,452
Order 0.61 1.03 0 0 0 1.00 430.00 123,748,452
UniqueItem 1.42 2.93 0 0 0 2.00 4,191.00 123,748,452
UniqueBrand 1.15 2.09 0 0 0 2.00 332.00 123,748,452
UniqueSubcategory 1.00 1.70 0 0 0 2.00 64.00 123,748,452
UniqueCategory 0.85 1.35 0 0 0 1.00 18.00 123,748,452
Transactions in Offline Channel
Spend 195.07 631.18 0 0 0 125.93 2,340,617.16 123,748,452
Quantity 1.78 9.03 0 0 0 1.00 44,158.00 123,748,452
Orders 0.43 0.87 0 0 0 1.00 430.00 123,748,452
UniqueItems 1.11 2.72 0 0 0 1.00 4,191.00 123,748,452
UniqueBrands 0.90 1.94 0 0 0 1.00 332.00 123,748,452
UniqueSubcategories 0.78 1.59 0 0 0 1.00 64.00 123,748,452
UniqueCategories 0.66 1.26 0 0 0 1.00 18.00 123,748,452
Transactions in Online Channel
Spend 107.95 421.24 0 0 0 0 276,399.66 123,748,452
Quantity 0.66 4.68 0 0 0 0 4,800.00 123,748,452
Orders 0.18 0.56 0 0 0 0 130.00 123,748,452
UniqueItems 0.32 1.14 0 0 0 0 141.00 123,748,452
UniqueBrands 0.27 0.88 0 0 0 0 44.00 123,748,452
UniqueSubcategories 0.24 0.75 0 0 0 0 22.00 123,748,452
UniqueCategories 0.22 0.64 0 0 0 0 13.00 123,748,452
We quantify consumers’ purchase behavior on three dimensions: Volume, Frequency, and Variety.
  • $\bullet$
    Volume of purchase is operationalized using two variables – (1) $(\textit{Spend}_{it})$, which is the revenue generated from the customer $i$ in month $t$, and (2) $(\textit{Quantity}_{it})$, the total number of items purchased by customer $i$ in month $t$.
  • $\bullet$
    Frequency is measured by the number of unique orders per month $(\textit{Orders}_{it})$. This metric focuses on purchase frequency without considering the value of purchases, which differs from volume metrics.
  • $\bullet$
@@ 3.2.2 Customer Monthly Purchase Behaviors
To preserve the confidentiality of the data provider, we mask the true spend value ($\textit{Spend}_{it}$) by multiplying it by an undisclosed scalar. All the descriptive and model estimation results are thus shown in terms of Masked Currency Units (MCU). The rest of the variables listed above are shown and used as is, without any masking.
Table 4 presents customer-month-level summary statistics of all variables discussed above for three cases: (1) Transactions across both channels, (2) Transactions in the offline channel, and (3) Transactions in the online channel. Table 4 shows the mean statistic of each variable, measured at the customer-month-channel level, which is calculated as $\frac{\sum_{i=1}^{N}\sum_{t=1}^{T_{i}}X_{itc}}{\sum_{i=1}^{N}T_{i}}$, where $X_{itc}$ is the customer-month-channel level variable, $T_{i}$ is the panel length of customer $i$, and $N$ is the total number of customers, which is 5,934,248. $\sum_{i=1}^{N}T_{i}$ is the total number of observations (referred to as count in Table 4), which is 123,748,452.
We find that, on average, consumers spend more offline than online. It is important to note that although the unconditional mean of monthly offline spend (195.07 MCU) is higher than online spend (107.95 MCU), we cannot conclude that, on average, customers spend more money offline than online. This is because some customers may only access one sales channel, or they strategically select the channel that benefits them the most, or choose different channels for different types of purchases. We will discuss these ideas in detail later in the paper. Further, we find that all the variety and frequency variables are also higher for offline purchases compared to online purchases. All the purchase metrics have over 50% of observations as zero, which suggests that purchase frequencies are lower than once per month for a large fraction of consumers, though some consumers purchase very frequently and buy a large variety of products.
4 Descriptive Analysis
As a first step, we conduct some descriptive analyses to examine (a) cross-sectional differences in demographics, monthly spending, and other purchase behaviors of multichannel consumers, online-only consumers, and offline-only consumers; and (b) within-customer changes to monthly spending and other purchase behaviors after adopting online shopping. We highlight the main takeaways from each analysis here and refer readers to Web Appendix B for additional details.
First, we observe that multichannel consumers are more likely to be female compared to customers who are online-only or offline-only. Of the three groups, online-only consumers are the youngest, and offline-only consumers are the oldest. Multichannel consumers have the highest income, and offline-only consumers have the lowest income. Next, we run simple cross-sectional regressions for the outcome variables, spend, quantity, unique items, brands and categories, on the explanatory variable of consumer-type (online-only, offline-only and multichannel), controlling for demographics and month fixed effects, to measure cross-sectional differences in spend and other purchase behaviors. We find that multichannel consumers spend more (and also purchase higher quantities, more unique items, brands, and categories) relative to both offline-only and online-only consumers. These patterns align with previous literature (Kumar and Venkatesan 2005; Thomas and Sullivan 2005; Kushwaha and Shankar 2013) documenting that multichannel consumers are more valuable than single-channel consumers (though the set of channels studied in that literature is different).
Next, we conduct a within-consumer analysis to examine whether consumers who were initially only shopping offline increase their spending after adopting online shopping. We do see a significant increase in monthly spending (by about 39%) post-adoption  [^5] Referring to Table A6 in Web Appendix B.2, this number is given by the incremental spend after online adoption (124.2 MCU) divided by the pre-adoption average spend (318.4 MCU). These results are consistent with earlier work (studying different channel transitions) that shows that consumers who transition to shopping on multiple channels from previously using only a single channel tend to spend more (Venkatesan and Ravishanker 2007; Ansari, Mela, and Neslin 2008) post multichannel adoption. We also see similar increases in quantity, unique items, brands, and categories.
While these descriptive analyses confirm existing studies, they nevertheless have several shortcomings. The cross-sectional analysis does not measure changes in behavior, while the within-customer analysis assumes that the before-after changes we observe are entirely attributable to the adoption of online shopping (since there is no control group of non-adopters). Moreover, these analyses do not shed light on whether the different reasons for the adoption of online shopping affect consumers’ purchase behavior in the post-adoption period differently. We address these issues and adopt a more comprehensive empirical framework to answer our research questions in the next two sections.
5 Firm’s Problem and Sample Construction
We start by formalizing the firm’s problem. As discussed in $\S$1, our first goal is to quantify how the adoption of digital shopping channels (e-commerce website and/or mobile app) affects the purchase behavior of consumers who were previously purely offline-only (brick-and-mortar) shoppers in retail markets. Specifically, we seek to document the impact of the adoption of online shopping on the following three metrics: (1) Total Spend, (2) Profitability, and (3) Stickiness to the offline channels post-adoption. Formally, we are interested in the Average Treatment Effect on Treated (ATT) of adopting online shopping on these three metrics. A second goal is to understand whether the different reasons for adopting online shopping affect consumers’ post-adoption purchase behavior, and to what extent. To that end, we consider four types of reasons behind online adoption: (1) a macro-environmental shock (COVID-19), (2) a short-term promotion (Black Friday sales), (3) a long-term loyalty program, and (4) organic adoption, not driven by any external events. The first event is outside the firm’s control, whereas the latter two were firm-driven marketing activities. For each event, we first compare consumers who adopted online shopping adopters with a control group of offline-only consumers to study the effect of online adoption. Next, we compare consumers who adopted online shopping due to the event (i.e., event adopters) with those who adopted organically (without external incentives) around the same time frame as the event (i.e., organic adopters) to understand how these different external events affect post-adoption behavior relative to organic adoption. In the process, we also seek to provide theoretical underpinnings for how and why the different reasons for online adoption affect post-adoption purchase behavior differently.

### v2

@@ 2 Related Literature
    Moreover, our setting is a fast-moving consumer packaged goods retailer (pet supplies), whereas much of the earlier literature examines books, apparel, video games, electronics, and technology. These differ from our context along several dimensions, including purchase frequency, competitive intensity, and product type (often more hedonic than utilitarian), which has been shown to matter for multichannel effects (Kushwaha and Shankar 2013).
Thus, we extend the multichannel literature by (i) examining how post-adoption behavior varies with channel adoption pathways and providing theoretical explanations for these differences, and (ii) focusing on a practically important channel transition—offline-only to combined offline and online usage—that has not been separately studied.
From a methodological perspective, prior research has used three broad approaches to documenting the value of adopting new channels: cross-sectional, within-customer, and difference-in-differences (DID). Early work (Thomas and Sullivan 2005; Kumar and Venkatesan 2005; Kushwaha and Shankar 2013; Montaguti, Neslin, and Valentini 2016) uses cross-sectional comparisons of multichannel versus single-channel customers and concludes that multichannel customers tend to be more valuable. However, cross-sectional comparisons lack individual-specific controls and cannot isolate how behavior changes after multichannel adoption. A second stream (Venkatesan and Ravishanker 2007; Ansari, Mela, and Neslin 2008) employs within-customer analysis, comparing a customer’s purchase behavior before versus after adopting multiple channels. This work also finds higher spending after multichannel adoption, but without an appropriate control group of non-adopters, it requires strong assumptions that no other time-varying factors affect behavior. More recent work (Wang, Malthouse, and Krishnamurthi 2015; Narang and Shankar 2019) uses DID with non-adopters as a control group, addressing many of these concerns. Our paper also employs a DID framework with non-adopters as a control group, but brings this methodology to a new setting and focuses on heterogeneity across adoption pathways, which has not been explored in prior empirical work.
3 Background, Goals, and Theory
As discussed in $\S$2, the existing multichannel literature has not considered heterogeneity in channel adoption pathways and has implicitly assumed that all multichannel customers behave similarly, regardless of why they adopted new channels to shop with a given retailer. However, consumers may adopt a retailer’s online shopping channels for many different reasons, and it is plausible that these different pathways affect post-adoption behavior. At a high level, there are three broad classes of online adoption pathways: (1) organic adoption, where customers voluntarily adopt a retailer’s online shopping channels due to the added convenience of an additional channel; (2) macro-environmental shocks, such as public health crises (e.g., COVID-19), nearby store closures, or severe weather events that prevent store visits; and (3) firm-driven promotions, where retailers explicitly incentivize adoption of their online channels.
Each pathway represents a different type of consumer self-selection, i.e., different types of consumers choose to adopt a retailer’s online shopping channels in response to each stimulus. Our goal is not to rule out selection effects; rather, we seek to document whether and how different reasons for adopting online shopping with a retailer affect post-adoption behavior. From a managerial perspective, firms cannot randomize online adoption across consumers; they can only “encourage” adoption through promotional activities and observe how different encouragements translate into different adoption pathways and subsequent purchases. [^2] We refer readers to the broader literature on “encouragement design” in marketing and the social sciences. Broadly, this literature distinguishes between interventions where firms can assign consumers to treatment arms (e.g., pricing, advertising, web page design) and those where consumers must self-select into the focal action (e.g., signing up for a newsletter, adopting online shopping) (Bradlow 1998; Syrgkanis et al. 2019; Mummalaneni, Yoganarasimhan, and Pathak 2023). In the latter, firms can only encourage the action (e.g., via promotions), and even a fully randomized experiment cannot eliminate selection because the outcome of interest (online adoption) is itself chosen by the consumer. In such settings, the natural estimand is the Average Treatment Effect on the Treated (ATT), rather than the Average Treatment Effect (ATE). Accordingly, we aim to document the ATT of online adoption on post-adoption behavior for different pathways and to relate differences in these ATTs to theories of consumer behavior.
We next discuss behavioral mechanisms that link channel adoption pathways (organic, macro shocks, and promotions) to post-adoption outcomes. First, customers who adopt online channels organically may behave differently from those who adopt due to external factors. Organic adoption most plausibly signals genuine preference or readiness to use online channels, rather than being forced or incentivized by external shocks or promotions. Consumer inertia and habit theory (Verplanken and Wood 2006; Wood and Neal 2009) suggest that adoption driven by external factors may temporarily disrupt existing routines, but absent reinforcement, consumers may revert toward prior habits and use the newly adopted channels less intensively. Organic adopters, by contrast, are more likely to invest in learning and to enjoy the convenience of online channels, leading to more sustained use.
If so, we can make several predictions about differences between organic adopters and externally driven adopters. (1) The post-adoption share of spending on online channels should be higher for organic adopters than for externally driven adopters. (2) Because profit margins typically differ across channels, this difference in online share may translate into lower (higher) profitability for organic adopters if online channels have lower (higher) margins than offline channels. Prior work also shows that older consumers tend to be more entrenched in their habits and exhibit higher inertia than younger consumers (Yang and Ching 2014). Thus, we expect divergence in online usage between organic adopters and externally driven adopters to be more pronounced when external factors disproportionately trigger adoption among older consumers (e.g., public health concerns like COVID-19 or promotions targeted at older segments). This leads to a third prediction: (3) the gap in online share between organic adopters and externally driven adopters should be larger when the external stimulus induces relatively older consumers to adopt.
Promotions offering discounts or rewards in return for increased spending during a limited period may also lead to distinctive post-adoption behavior relative to other external shocks. Prior research documents forward buying and demand distortion (Neslin, Henderson, and Quelch 1985; Hendel and Nevo 2006; Van Heerde and Neslin 2017), whereby time-limited promotions induce stockpiling and reduce demand in the post-promotion period. This mechanism is particularly relevant for promotions designed to induce customers to start shopping online and motivates a fourth prediction: (4) consumers who adopt in response to promotions may increase spending during the promotional period but reduce spending afterwards, thereby attenuating (or even reversing) the longer-run spending boost normally associated with multichannel adoption.
Based on these mechanisms, we seek to empirically test the following predictions:
  • $\bullet$
    H1: Customers who adopt a firm’s online channels and become multichannel shoppers will increase their spending post-adoption relative to offline-only customers (consistent with prior findings on multichannel shoppers; see $\S$2).
  • $\bullet$
    H2: Customers with different channel adoption pathways will exhibit different post-adoption behavior in terms of spend, profitability, and channel mix. These differences will arise from a combination of customer demographics, consumer habit theory, and forward buying. Specifically,
    • $\bullet$
      H2a: Consumers who adopt in response to time-limited promotions may increase spending during the promotional period but reduce spending thereafter, dampening (or reversing) the spending boost from multichannel adoption in the post-promotion period.
    • $\bullet$
      H2b: Externally driven adopters (vs. organic adopters) may be less likely to show the same level and growth in the share of spending on the newly adopted online channel, as predicted by consumer inertia and habit theory.
    • $\bullet$
      H2c: The difference in online share between organic adopters and externally driven adopters may be larger when external factors cause relatively older consumers to adopt.
    • $\bullet$
      H2d: Differences in online share may translate into lower (higher) profitability for organic adopters if online channels have lower (higher) margins relative to offline channels.
@@ 4.1 Setting
Recall that the key goals of the paper are: (1) to empirically document the effect of adopting online shopping on post-adoption consumption behavior, (2) to test whether different channel adoption pathways lead to different post-adoption behaviors, and (3) to verify that the differences in post-adoption behavior can be predicted using theoretical underpinnings of purchase behavior in retail settings, so as to provide generalizable insights for managers and researchers. To accomplish these goals, we need data from a setting where we can observe not only when an offline-only consumer adopts online shopping but also the driver of this decision, i.e., the adoption pathway. Further, we need sufficient variants in channel adoption pathways to meaningfully connect the post-adoption behavior to theory. Therefore, we focus on a retail data source that satisfies these requirements.
Our data comes from one of the largest Brazilian pet supplies retailers, which sells a wide variety of products for pets, such as pet food and snacks, medicines/pharmacy items, hygiene and grooming products, and pet toys and accessories. Pet supplies is a large and competitive industry with about \$250 billion in annual revenue globally (Fortune 2024). The firm has both offline and online channels. The retailer operates more than 200 physical stores in Brazil, which serve as the offline channel, and has a website as well as a mobile app, which serve as the online channel. Moreover, our data span 2019-2024 during which there were three different events including a macro/environmental shock (COVID-19) and two different types of promotions designed to incentivize online shopping (discussed in more detail in $\S$5) which led to many consumers adopting online shopping with the firm. Further, our data are quite rich, and we are able to observe when a consumer adopts online shopping (and thus can connect the timing of adoption to the channel adoption pathway) and both the pre- and post-adoption consumption behavior. For these reasons, it is an ideal setting for our study.
@@ 4.2 Data
Our data includes transaction-level records of all individual customers from both online and offline channels during the 63-month period from January 1st, 2019, to March 31st, 2024. Each transaction record is uniquely identified by an order number with one or more purchased items. For each purchased item, we have information on the unit price, quantity purchased, brand, purchased channel (online/offline), subcategory, and category (e.g., food, hygiene).
@@
5 Estimation Sample and Summary Statistics